According to Masterfoods, the company that manufactures M&Ms, 12% of peanut M&Ms are brown, 15% are yellow, 12% are red, 23% are blue, 23% are orange and 15% are green. Assume that selecting multiple M&Ms are independent events. (Round your answers to three decimal places, for example 0.123)
Compute the probability that a randomly selected peanut M&M is not green.

Compute the probability that three randomly selected peanut M&Ms are all green.

If you randomly select four peanut M&Ms, compute the probability that none of them are red.

Answer by Boreal(15235)

p(not green) is 1-p(green)=0.85
-
all three are green is 0.15^3=0.003375 or 0.003. Sorry, I was still thinking of complement.
-
none red is 0.88^4=0.5997 or 0.600.
There is an 88% chance it won't be red.